OpenStudy (anonymous):

Solve for a. a/4+a/3=21 3 36 147

OpenStudy (anonymous):

\[3a/12+4a/12=21\]\[7a/12=21\]\[a/12=3\] \[a=36\] 36

OpenStudy (auctoratrox):

36